please give an example to solve logs.....

Grade:8

1 Answers

Sunil Kumar FP
9 years ago
remember to solve problem of log folowing formula sould be kept in mind
(1)loga (b) =loga/logb
(2)log a (b) =k,then a^k=b,
(3)log a -log b =log(a/b)
(4)log a+ log b =log(ab)
(5)log a =b ,then a=antilog (b)
These formulas should be kept in mind to solve the problem in logarithm .These are more than enough to solve any problem in logarithm.
For any further doubt please ask here we will be solving the numerical